# PLAICraft: Large-Scale Time-Aligned Vision-Speech-Action Dataset for Embodied AI

### v1

@@ A.8 Content Moderation
A.9 Data Pre-processing
@@ A.9.4 Data Encoding
Audio signals, encompassing both and , are encoded utilizing the EnCodec 24 kHz model. Following prior approaches [Shen et al., 2023, Evans et al., 2024], we employ continuous latent vectors produced by this model, characterized by dimensions with a frequency of 75Hz.
@@ A.11 Data Versioning
A.12 DataLoader and Batch Sampler Design.
Our dataloader functions with a global metadata database, which contains game sessions metadata like who played the session, what modalities are available in these recordings, the start time of each gameplay session, etc. The database download is also available at https://blog.plaicraft.ai/2025/05/11/download-user-data/. The dataloader can then load the data from any specified sessions or players or from the entire database. As mentioned previously, due to the frequency discrepancy between video and audio, we treat 2 video latent frames, i.e. a 200ms window, as our minimal unit. Thus the dataloader will return a dict object of the following content:
| Metadata: | |||
| Video tensor shape: | torch.Size([B, T, 2, 4, 96, 160]) | ||
| audio_speak shape: | torch.Size([B, T, 15, 128]) | ||
| audio_hear shape: | torch.Size([B, T, 15, 128]) | ||
| Mouse movement shape: | torch.Size([B, T, 2, 10, 2]) | ||
| Key press shape: | torch.Size([B, T, 2, 5, 16]) | ||
@@
Appendix B Evaluation Tasks
B.1 Complete Evaluation Prompt Table
@@
| Test Type | Prompt | Response | Metric | CHC | Video | Pstart (ms) | Pend (ms) | Rstart (ms) | Rend (ms) |
| audio to action binding | Please press Space. | Space pressed. | 1 | 65184 | 67045 | 67970 | 68120 | ||
| audio to action binding | Please press Shift. | Shift pressed. | 1 | 70366 | 71666 | 73120 | 73650 | ||
| audio to action binding | Please press T. | T pressed. | 1 | 86852 | 87812 | 89240 | 89410 | ||
@@
| spelling ability | Please press the letter that ends the word STOP, as in S-T-O… What letter was that? | P pressed. | 1 | 104752 | 113036 | 115950 | 116070 | ||
@@
| spelling ability | Please press the letter that ends big, as in B-I. What letter did you press? Did you press the letter G? | G pressed. | 1 | 117057 | 117697 | 119133 | 119373 | ||
| audio to action binding | Please press the letter Q. | Q pressed. | 1 | 121790 | 122531 | 123195 | 123515 | ||
| action recognition | What did you just close? | Door. | 1 | 145813 | 146594 | 151430 | 152860 | ||
| internal state recognition | How much? What is that green? What is that green dot? | XP. | 1 | 474250 | 475731 | 476517 | 476837 | ||
@@
| world model understanding | I’m going to shoot you with an arrow, okay? And I want you to tell me how many hearts of damage are done, okay? | Two. | 1 | 508154 | 508954 | 510590 | 511010 | ||
@@
| object recognition | What do you have in your right hand? | Carrot. | 1 | 501410 | 502551 | 548097 | 548557 | ||
@@
| object recognition | What do you have in your left hand? | Torch. | 1 | 554727 | 557228 | 559065 | 559325 | ||
| color understanding | What color sheep am I standing next to? | Black. | 1 | 545784 | 547284 | 567047 | 567367 | ||
| color understanding | What color sheep am I standing next to? | Gray. | 1 | 545784 | 547284 | 548097 | 548557 | ||
@@
| color understanding | What color flower am I standing next to and looking at? | White. | 1 | 554727 | 557228 | 559065 | 559325 | ||
| color understanding | What color of flower am I looking at? | Red. | 1 | 564107 | 565929 | 567047 | 567367 | ||
| color understanding | What color of flower am I looking at? | Yellow. | 1 | 564107 | 565929 | 572471 | 572892 | ||
@@
| type recognition | What is the stack of stuff that I’m standing next to made out of? | wool. | 1 | 862534 | 863555 | 1111207 | 1111607 | ||
@@
| arithmetic | If you add up the number of blocks in the wool pile and the number of blocks in the granite pile, how many blocks total are there? | Five. | 1 | 1142820 | 1144281 | 1146213 | 1146573 | ||
@@
| arithmetic | Now, if you add up the number of blocks in the granite pile and the wool pile, how many blocks are there total? | Four. | 1 | 1180221 | 1185165 | 1243900 | 1244180 | ||
@@
| arithmetic | I am going to place a block in the granite pile. How many blocks are in the granite pile now? | Four. | 1 | 1233210 | 1234591 | 1243900 | 1244180 | ||
| object constancy | Now, jump up. Can you see me? | yes. | 2 | 602432 | 604614 | 605865 | 606185 | ||
| object constancy | Jump up. Did I move? | no. | 2 | 613679 | 614439 | 615528 | 615788 | ||
| audio to action binding | Press W. | W pressed. | 2 | 677774 | 678475 | 679650 | 679740 | ||
| audio to action binding | Press S. | S pressed. | 2 | 682398 | 682998 | 684010 | 684390 | ||
| audio to action binding | Press A. | A pressed. | 2 | 687001 | 687641 | 689700 | 689850 | ||
| audio to action binding | Press D. | D pressed. | 2 | 692265 | 692625 | 692750 | 692940 | ||
| audio to action binding | Press S. | S pressed. | 2 | 682398 | 682998 | 697070 | 697150 | ||
| audio to action binding | Do it again. Press S. | S pressed. | 2 | 707405 | 708345 | 709500 | 709700 | ||
| audio to action binding | Press W. | W pressed. | 2 | 677774 | 678475 | 713210 | 721000 | ||
@@
| working memory, object constancy | Have you been here before? | Yes. | 2 | 1452930 | 1453691 | 1482922 | 1483343 | ||
@@
| working memory, object constancy | Will we be someplace we’ve been before or will we be someplace new? So say old or new. | old. | 2 | 1786469 | 1792373 | 1797021 | 1797241 | ||
| object recognition | Can you see a spider? | Yes. | 2 | 1850053 | 1850994 | 223010 | 223250 | ||
| object recognition | Can you see any sheep? | Yes. | 2 | 1854695 | 1855656 | 223010 | 223250 | ||
| type recognition | What kind of block am I standing on? | Sand. | 2 | 1923506 | 1924907 | 1926608 | 1926968 | ||
| object recognition | What am I in right now? | Water. | 2 | 1938866 | 1939747 | 1940424 | 1940731 | ||
| world model understanding | Is the sheep dead? | no. | 3 | 101422 | 102222 | 103176 | 103456 | ||
| world model understanding | Is the sheep dead? | yes. | 3 | 101422 | 102222 | 110801 | 111221 | ||
| symbol recognition, numerosity | how many pieces of wool are in the box? | Four. | 3 | 316272 | 318352 | 319976 | 320256 | ||
| symbol recognition, numerosity | How many different kinds of things are in the box? | Two. | 3 | 322233 | 324093 | 324998 | 325158 | ||
| symbol recognition, numerosity | How many pieces of wood are in the box? | six. | 3 | 291773 | 293074 | 369764 | 370124 | ||
| type recognition | What is the first thing that I put in the box? | Wood. | 3 | 483807 | 485687 | 488025 | 488445 | ||
| spatial understanding | Approximately how many blocks away is the box? | hundred. | 3 | 578622 | 580584 | 583714 | 584054 | ||
| object recognition | Morgan, in front of me, what is that? | zombie. | 3 | 846473 | 848094 | 849320 | 849741 | ||
| object recognition | Morgan, are those torches? | Yeah. | 3 | 891198 | 892679 | 895606 | 896186 | ||
| symbol recognition, numerosity | How many pieces of wood? | Six. | 3 | 1668117 | 1669038 | 1670080 | 1670321 | ||
@@
| Reading | What is the name of the last person who left the game? | Brent. | 3 | 1902286 | 1904908 | 1908101 | 1908621 | ||
@@
| Reading | Okay, what is the name of the person who joined just after me? | Chloe. | 3 | 1936849 | 1946756 | 1948190 | 1948630 | ||
@@
| long term memory | What is the middle item in the chest? | Carrot. | 3 | 2171282 | 2172942 | 2174829 | 2175269 | ||
B.2 Evaluation Prompt Visual Examples
| Frame Prompt | Prompt | Response |
How many sheep do you see? | Two. |
Can you see the sun? | Yes. |
Can you see the sun now? | No. |
Am I facing the direction of the box? | No. |
Am I facing the direction of the box? | Yes. |
| Frame Prompt | Prompt | Response |
What animal do you see? | Sheep. |
@@ B.2 Evaluation Prompt Visual Examples
Jump one time. Can you see me? | No. |
Jump one time. Can you see me? | Yes. |

### v2

@@ A.5 Content Moderation
A.6 Recording Software
Video & Audio Video recordings are captured using Open Broadcaster Software (OBS) running on the EC2 instance. Videos are recorded at a constant frame rate of 30 FPS with a resolution of 1280x720, encoded using the H.264 codec. OBS is specifically configured to capture only the Minecraft application window in full-screen mode, ensuring that all recorded frames pertain solely to gameplay. Audio streams are simultaneously recorded by OBS, separately capturing user microphone input and game audio output. Each audio stream is recorded at a constant sampling rate of 48 kHz in stereo format, encoded with the AAC codec. We take the exact timestamp when OBS’s ffmpeg muxer began writing the file as the start timestamp, and align all modalities to it.
Mouse & Keyboard Mouse and keyboard inputs are precisely captured using a custom-developed Forge mod integrated directly into the client-side Minecraft application. This method provides superior timing accuracy by recording when the game server actually processes input events, unlike traditional system-level logs. Input events for both mouse and keyboard are recorded on an event-driven basis with millisecond precision, maintaining a maximum polling rate of 100 Hz, resulting in a minimum interval of 10 ms between consecutive data points. Additionally, comprehensive system-level logging using pyxhook is implemented as a secondary data collection measure. These system-level logs, although less precise in timestamp accuracy, serve primarily to validate the integrity and completeness of the mod-generated input data.
A.7 Data Pre-processing
@@ A.7.4 Data Encoding
Audio signals, encompassing both and , are encoded utilizing the EnCodec 24 kHz model. Following prior approaches [26, 13], we employ continuous latent vectors produced by this model, characterized by dimensions with a frequency of 75Hz.
@@ A.9 Data Versioning
A.10 Data Unit Format
Our dataloader functions with a global metadata database, which contains game sessions metadata like who played the session, what modalities are available in these recordings, the start time of each gameplay session, etc. The dataloader can then load the data from any specified sessions or players or from the entire database. As mentioned previously, due to the frequency discrepancy between video and audio, we treat 2 video latent frames, i.e. a 200ms window, as our minimal unit. Thus the dataloader will return a dict object as illustrated in Table˜8.
| Key | Shape |
| Metadata | [B] |
| Video tensor | [B, T, 2, 4, 96, 160] |
| audio_speak | [B, T, 15, 128] |
| audio_hear | [B, T, 15, 128] |
| Mouse movement | [B, T, 2, 10, 2] |
| Key press | [B, T, 2, 5, 16] |
@@ A.10 Data Unit Format
A.11 Automatic Data Annotation
To identify subsets of highly interactive data, particularly relevant for Embodied AI, we developed a methodology leveraging audio transcribed by WhisperX [3], and the capabilities of large language models (LLMs). We first defined highly interactive data as continuous, game-related interactions between at least two players, within a specific time window. Notably, a key indicator of interactivity was the presence of rich, conversational audio. We used WhisperX-generated audio transcripts to create a dataset and identify these interactions. To preserve conversational coherence, we split the transcript into segments such that a new segment begins after a pause of at least 10 seconds between utterances. These segments represent potential areas of interactivity and Minecraft social interactions. Given the scale of the resulting dataset, manual annotation of segments was infeasible. Instead, we leveraged recent advances in large language models to perform automated annotation. A binary classification approach using a 4-shot prompt demonstrated high accuracy and efficiency on a validation subset. Based on this evaluation and overall cost considerations, we used models such as DeepSeek-V3 [10] and DeepSeek-R1 [9] to label the full dataset with binary annotations. Table 4 demonstrates the labeling scheme. Segments assigned specific labels were then identified as highly interactive and selected for further analysis. Figure˜10 showcases the interaction portion of a few randomly-selected players (the names are the in-game assigned names) and the weekly accumulation of their highly interactive hours.
A.12 Additional Visual Examples
We demonstrate more visual examples in Fig.˜13, Fig.˜14, Fig.˜15. The visualization follows the same format as Fig.˜2, but here we also depict temporal progression: frames within each row correspond to a single example, are spaced 1 second apart, and progress from left to right in time.
A.13 Complete Evaluation Prompt Table
@@
| ID | Test Type | Spoken Instruction | Response | Metric | CHC | Video | Rstart (ms) | Rend (ms) |
| 1 | audio to action binding | Please press Space. | Space pressed. | 1 | 67970 | 68120 | ||
| 2 | audio to action binding | Please press Shift. | Shift pressed. | 1 | 73120 | 73650 | ||
| 3 | audio to action binding | Please press T. | T pressed. | 1 | 89240 | 89410 | ||
@@
| 4 | spelling ability | Please press the letter that ends the word STOP, as in S-T-O… | P pressed. | 1 | 115950 | 116070 | ||
@@
| 5 | spelling ability | Please press the letter that ends big, as in B-I… | G pressed. | 1 | 136930 | 137050 | ||
| 6 | audio to action binding | Please press the letter Q. | Q pressed. | 1 | 161850 | 162020 | ||
| 7 | action recognition | What did you just close? | Door. | 1 | 200116 | 200550 | ||
| 8 | internal state recognition | What is that green dot? | XP. | 1 | 435183 | 435866 | ||
@@
| 9 | world model understanding | I’m going to shoot you with an arrow, okay? And I want you to tell me how many hearts of damage are done, okay? | Two. | 1 | 482750 | 483500 | ||
@@
| 10 | object recognition | What do you have in your right hand? | Carrot. | 1 | 510550 | 511016 | ||
@@
| 11 | object recognition | What do you have in your left hand? | Torch. | 1 | 514683 | 515383 | ||
| 12 | color understanding | What color sheep am I standing next to? | Black. | 1 | 534633 | 535216 | ||
| 13 | color understanding | What color sheep am I standing next to? | Gray. | 1 | 548100 | 548566 | ||
@@
| 14 | color understanding | What color flower am I standing next to and looking at? | White. | 1 | 558866 | 559500 | ||
| 15 | color understanding | What color of flower am I looking at? | Red. | 1 | 566966 | 567433 | ||
| 16 | color understanding | What color of flower am I looking at? | Yellow. | 1 | 572383 | 572899 | ||
@@
| 17 | type recognition | What is the stack of stuff that I’m standing next to made out of? | Wool. | 1 | 1111099 | 1111616 | ||
@@
| 18 | arithmetic | If you add up the number of blocks in the wool pile and the number of blocks in the granite pile, how many blocks total are there? | Five. | 1 | 1182683 | 1183299 | ||
@@
| 19 | arithmetic | Now, if you add up the number of blocks in the granite pile and the wool pile, how many blocks are there total? | Four. | 1 | 1199099 | 1199599 | ||
@@
| 20 | arithmetic | I am going to place a block in the granite pile. How many blocks are in the granite pile now? | Three. | 1 | 1235883 | 1236333 | ||
| 21 | object constancy | Now, jump up. Can you see me? | Yes. | 2 | 605866 | 606200 | ||
| 22 | object constancy | Jump up. Did I move? | No. | 2 | 615450 | 615899 | ||
| 23 | audio to action binding | Press W. | W pressed. | 2 | 679650 | 679740 | ||
| 24 | audio to action binding | Press S. | S pressed. | 2 | 684010 | 684390 |
| ID | Test Type | Spoken Instruction | Response | Metric | CHC | Video | Rstart (ms) | Rend (ms) |
| 25 | audio to action binding | Press A. | A pressed. | 2 | 689700 | 689850 | ||
| 26 | audio to action binding | Press D. | D pressed. | 2 | 692750 | 692940 | ||
| 27 | audio to action binding | Press S. | S pressed. | 2 | 697070 | 697150 | ||
| 28 | audio to action binding | Do it again. Press S. | S pressed. | 2 | 709500 | 709700 | ||
| 29 | audio to action binding | Press W. | W pressed. | 2 | 713210 | 713440 | ||
@@
| 30 | working memory, object constancy | Have you been here before? | Yes. | 2 | 1456133 | 1456466 | ||
@@
| 31 | working memory, object constancy | Will we be someplace we’ve been before or will we be someplace new? So say old or new. | Old. | 2 | 1796900 | 1797250 | ||
| 32 | object recognition | Can you see a spider? | Yes. | 2 | 1852349 | 1852683 | ||
| 33 | object recognition | Can you see any sheep? | Yes. | 2 | 1857016 | 1857549 | ||
| 34 | type recognition | What kind of block am I standing on? | Sand. | 2 | 1926616 | 1927033 | ||
| 35 | object recognition | What am I in right now? | Water. | 2 | 1940433 | 1940783 | ||
| 36 | world model understanding | Is the sheep dead? | No. | 3 | 103183 | 103466 | ||
| 37 | world model understanding | Is the sheep dead? | Yes. | 3 | 110816 | 111266 | ||
| 38 | symbol recognition, numerosity | How many pieces of wool are in the box? | Four. | 3 | 319983 | 320266 | ||
| 39 | symbol recognition, numerosity | How many different kinds of things are in the box? | Two. | 3 | 324983 | 325166 | ||
| 40 | symbol recognition, numerosity | How many pieces of wood are in the box? | Twelve. | 3 | 346466 | 346983 | ||
| 41 | type recognition | What is the first thing that I put in the box? | Wood. | 3 | 487899 | 488516 | ||
| 42 | spatial understanding | Approximately how many blocks away is the box? | Hundred. | 3 | 583549 | 584283 | ||
| 43 | object recognition | Morgan, in front of me, what is that? | Zombie. | 3 | 849333 | 849750 | ||
| 44 | object recognition | Morgan, are those torches? | Yeah. | 3 | 895549 | 896200 | ||
| 45 | symbol recognition, numerosity | How many pieces of wood? | Six. | 3 | 1669933 | 1670616 | ||
@@
| 46 | reading | What is the name of the last person who left the game? | Brent. | 3 | 1908099 | 1908866 | ||
@@
| 47 | reading | Okay, what is the name of the person who joined just after me? | Chloe. | 3 | 1942033 | 1942466 | ||
@@
| 48 | long term memory | What is the middle item in the chest? | Carrot. | 3 | 2174783 | 2175283 |
A.14 Evaluation Prompt Visual Examples
We pick a few prompts and show the nearest frames before response for visualization in Table 11 and 12.
| Nearest Frame | Spoken Instruction | Response |
@@ A.14 Evaluation Prompt Visual Examples
What do you have in your left hand? | Torch. |
@@
What color of flower am I looking at? | Red. |
What is the stack of stuff that I’m standing next to made out of? | Wool. |
| Nearest Frame | Spoken Instruction | Response |
I am going to place a block in the granite pile. How many blocks are in the granite pile now? | Three. |
Now, jump up. Can you see me? | Yes. |
Can you see a spider? | Yes. |
Can you see any sheep? | Yes. |
What am I in right now? | Water. |